Engberg Company installs lawn sod in home yards. The companys most recent monthly contribution format income statement follows: Percent of Sales 100% 40% Sales variable expenses Contribution margin Fixed expenses Net operating income Amount $80, 880 32,68 48, 000 38, 833 Required: 1. What is the companys degree of operating leverage? 2. Using the degree of operating leverage, estimate the impact on net operating income of a 5% increase in sales. 3, construct a new contribution format income statement ror the company assuming a 5% increase in sales. Complete this question by entering your answers In the tabs below Required 1 Required 2 Required 3 What is the companys degree of operating leverage? (Round your answer to 1 decimal place.) 4.8 Required 2 >
Answer #1

1) Degree of operating leverage = Contribution margin/EBIT = 48000/10000 = 4.8 Times

2) Net operating income increase by 4.8*5 = 24%

3) Contribution format income statement

Sales (80000*1.05) 84000
variable cost (32000*1.05) 33600
Contribution margin 50400
Fixed cost 38000
Net operating income 12400